Johnb:

The equation below shows the area of a trapezoid, A, with a height of 9 cm, and one base 35 cm: A = 9 over 2 (b + 35) Which of the following formulas correctly solves for the other base, b? (1 point) a b = 2A over 9 + 35 b b = 2A over 9 − 35 c b = 2 multiplied by A + 35, all over 9 d b = 2 multiplied by A â ' 35, all over 9

Angle:

Hello @Johnb and Welcome to QuestionCove! It looks like this is your equation \(A = \frac{9}{2}(b+35)\) Your goal is to try to undo the multiplication and undo the addition to try to get the ` b ` value by itself.

Angle:

I would suggest multiplying both sides by 2/9 then subtracting both sides by 35 if you are confused by what I mean by this, just let me know

Angle:

` I would suggest multiplying both sides by 2/9 ` \(\frac{2}{9}*A = \frac{2}{9}*\frac{9}{2}(b + 35)\) ^ try simplifying this

Angle:

` then subtracting both sides by 35 ` \(\frac{2}{9}*A ~~- 35=b ~+~35 ~-35\) simplifying this will get you your answer